# Universal Entanglement Revival of Topological Origin

### v1

@@ 3.2 Entanglement measures
Quantifying quantum entanglement [48] is vital in quantum information theory, yet is highly non-trivial. Common entanglement measures include entanglement entropy [49], concurrence [50], and logarithmic negativity (or the closely related negativity) [51, 52]. However, entanglement entropy is applicable only to pure states, while concurrence is practical only for two-qubit systems, whether in pure or mixed states. For mixed states of more general systems, the logarithmic negativity (or negativity) has been widely used as one of very few practically computable entanglement measures. Nevertheless, as it requires finding the eigenvalues of the density matrix after partial transposition, its computational cost increases quickly with system size. To avoid complications arising from partial transposition, quantum mutual information is often used as well. Although mutual information is not an entanglement measure in the strict mathematical sense, it still captures many important aspects of quantum correlations, as well as classical correlations. Below, we will investigate the entanglement content of the mixed state (14) using the logarithmic negativity. However, as we illustrate in A, the universal features we find are observed in the mutual information as well.
The logarithmic negativity between two subsystems $A$ and $B$ is defined as [51]
$\mathcal{E}(A,B):=\log_{2}\left\|\hat{\rho}^{T_{A}}\right\|=\log_{2}\left\|\hat{\rho}^{T_{B}}\right\|,$ (20)
@@
Recently, a new type of logarithmic negativity based on a partial time-reversal transformation has been suggested for fermionic systems [54, 55, 56]. Compared to the conventional (or bosonic) logarithmic negativity, this fermionic logarithmic negativity was shown to better capture the contributions to entanglement from Majorana fermions. However, the key advantage from our point of view lies in the fact that the partial time-reversal transformation of a fermionic Gaussian state remains Gaussian, allowing the fermionic logarithmic negativity to be calculated using a $2L\times 2L$ matrix of generalized Green’s functions [54]; that is, with a polynomial amount of resources and time. In our case, as the pairing correlations are zero, the fermionic logarithmic negativity is obtained directly from the $L\times L$ matrix of Green’s functions, see Eq. (19), and we could study half-SSH chains up to $L\sim 10^{3}$ without difficulties. For the bosonic case, the calculations are limited to relatively short chains.
3.3 Fermionic vs bosonic SSH models
Both bosonic and fermionic systems in the half-SSH model exhibit a chiral symmetry, characterized by the existence of a unitary operator that anti-commutes with the Hamiltonian. This symmetry leads to a degenerate zero-energy subspace and a symmetric single-particle energy spectrum around zero energy. Additionally, the edge mode at zero energy reflects the non-trivial topological character of the band structure.
For large bosonic systems, the size of the Hilbert space poses significant computational challenges. Certain techniques, such as block diagonalization of $\hat{\rho}^{T_{A}}$, are employed to alleviate the computational burden. In contrast, fermionic systems can benefit from the fact that the relevant state is Gaussian, since the partial time-reversal transformation of Eq. (16) remains Gaussian.
In the next Section, we will present our results for both bosonic and fermionic systems in parallel, to highlight the common universal behavior despite the differences summarized above. We attribute the common features of entanglement dynamics to the chiral symmetry and topological properties, especially the zero-energy edge states.
4 Results
We now present our findings, which reveal intriguing dynamical phenomena and provide deeper insights into the relationship between dynamics and topological properties. These results contribute to a better understanding of how topological effects influence dynamical processes, and have significant implications for future developments of quantum technologies.
Figure 3: Universal entanglement revival, obtained by computing the logarithmic negativity ($\mathcal{E}(\{2\},\{0,1,3,\dots\})$ between the single-site subsystem $\{\hat{f}_{2}\}$ and the rest of the chain. Panels (a) and (b) consider the fermionic model at $g_{1}/g_{2}=0.1$ and $0.5$, respectively. Panels (c) and (d) are for the bosonic model, also using $g_{1}/g_{2}=0.1$ and 0.5. Other parameters are $g_{2}=1$ (arbitrary units) and $\Gamma/g_{2}=0.1$.
4.1 Universal entanglement revival
Figure 3 summarizes our main results, concerning the evolution of entanglement between the second even site and the rest of the system. The two left panels of Fig. 3 refer to the fermionic half-SSH chain, where we computed the fermionic logarithmic negativity for two values of $g_{1}$ deep in the topologically non-trivial phase ($g_{1}\ll g_{2}$). The two right panels of Fig. 3 show the conventional logarithmic negativity of the bosonic model, using the same set of parameters. In the fermionic case, we could calculate the fermionic logarithmic negativity for systems as large as $L\sim 10^{3}$, and three representative sizes are shown in panels (a) and (b) of Fig. 3. In the bosonic case, the calculations were limited to relatively small values of $L$, because of the computational difficulties discussed in Section 3, and panels (c) and (d) show the results for $L=3,5,\dots,11$. In both cases, two distinctive features are clearly seen: First, the entanglement revives after an initial decrease. Considering the very fragile nature of quantum entanglement in the presence of decoherence, an entanglement revival is unusual and has attracted much interest [16, 1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 15, 14, 13]. Second, the temporal profile of the entanglement has a universal dependence, i.e., it is almost unaffected by the system size. Although in Fig. 3 we only show the entanglement between site $\hat{f}_{2}$ and the remaining sites, $\{\hat{f}_{0},\hat{f}_{1},\hat{f}_{3},\dots\hat{f}_{L}\}$, the presence of an universal revival behavior can be observed regardless of the specific choice of partitioning, as long as one subsystem is confined to the left edge. However, the visibility of the revival depends on the partitioning.
As demonstrated in a recent work [16], the entanglement revival behavior is due to a self-purification process directly related to the chiral symmetry and the associated zero-energy mode (not necessarily localized to edges): The system starts from a pure state with relatively large entanglement and, as it undergoes the decoherence process, decays to various intermediate states. Consequently, the purity of the state decreases significantly, which naturally leads to a considerable reduction in entanglement. After this intermediate stage, however, the system begins to recover its purity (and hence entanglement), because all particles have the fate of decaying out, except for the ones occupying the zero-energy edge mode. Eventually, the system stabilizes to the final state, i.e., the zero-energy mode, which is immune to the decay process. The entanglement in the steady state is always determined by this zero-energy mode.
While the above arguments explains the presence of the revival, its universal character is more intriguing, and is the main topic of the rest of the paper. In the following, we perform a detailed analysis of the size dependence, revealing that the universal behavior can be attributed to the topological feature of the system, especially the zero-energy edge mode.
4.2 Topological effects through the edge state
In Fig. 3, we have seen that the entanglement revival in the deep topological phase ($g_{1}/g_{2}\ll 1$) is universal, having a dynamical profile independent of system size. However, as the ratio $g_{1}/g_{2}$ increases to unity (i.e., the phase boundary of the topological phase is approached), the evolution of entanglement shows significant variations at small system sizes. When $g_{1}/g_{2}=0.9$, for example, the dynamical profile for $L=3$ is distinct from the profiles of larger systems ($L=19,1001$), as shown in the inset of Fig. 4. We find that the dynamical profiles remain universal as long as the system size is sufficiently large compared to a certain length scale, directly related to the localization length of the zero-energy edge mode. In the following, we examine this crossover scale in detail.
We define two length scales concerning the entanglement revival dynamics: The crossover length $L_{c}$ is defined as the minimum system size which, at a given ratio $g_{1}/g_{2}$, achieves a universal time dependence. In other words, the dynamical profiles of systems larger than $L_{c}$ are all universal (within a certain numerical tolerance $\epsilon$). More specifically, let $\mathcal{E}_{L}(t)$ be the logarithmic negativity for a system of length $L$. Then, choosing a sufficiently large $T$, at which the entanglement has reached its asymptotic value, we require that for all $L\geq L_{c}$
@@ 4.2 Topological effects through the edge state
Numerically, we approximate the $L\to\infty$ limit by taking $L=1001$. We also choose $T=20$ and $\epsilon=0.05$. Figure 4 plots the crossover length $L_{c}$ (blue empty circles) as a function of $g_{1}/g_{2}$. It is important to note that $L_{c}$ closely resembles the localization length $\xi$ of the edge mode (black solid line). For instance, when $g_{1}/g_{2}=0.9$, the crossover length matches the localization length $L_{c}\approx\xi\approx 19$. This value is highlighted by a horizontal dashed line in the main panel of Fig. 4. In the inset, we can see that the detailed entanglement profile with $L=19$ follows quite faithfully the universal dynamics at $L=1001$.
Figure 4: Comparison of the crossover length $L_{c}$ (blue empty circles) and the effective length $L_{\mathrm{eff}}$ (red filled circles) with the localization length $\xi$ of the edge state (black solid line), as a function of the coupling ratio $g_{1}/g_{2}$. The red dashed line marks the values $L_{c}\approx 19$, computed at $g_{1}/g_{2}=0.9$. Inset: time dependence of the logarithmic negativity between the single-site subsystem $\{\hat{f}_{2}\}$ and the rest of the chain. Here we take $g_{1}/g_{2}=0.9$ and $L=3$, $19$, $1001$ (from top to bottom).
@@
The explicit dependence of $L_{\rm eff}$ on $g_{1}/g_{2}$ is shown in Fig. 4. As seen, also the effective length $L_{\mathrm{eff}}$ (red filled circles) agrees well with the localization length $\xi$ of the edge mode. That both $L_{c}$ and $L_{\rm eff}$ follow closely the localization length $\xi$ strongly suggest that the spatial range of the edge mode is a critical factor in determining the onset of the universal entanglement revival behavior.
Another feature that supports the topological origin of the entanglement revival is the visibility, defined as the difference between the asymptotic value ($t\to\infty$) and the minimum of the logarithmic negativity. The details of the visibility differ between the fermionic and bosonic cases, and they are also affected by system parameters. However, as illustrated in Fig. 5 for a long fermionic chain, we find the following generic qualitative features: The visibility is zero at $g_{1}/g_{2}=0$ and first increases with $g_{1}/g_{2}$, reaching a maximum around $g_{1}/g_{2}\approx 0.5$. Then, it sharply drops to zero (i.e., the entanglement revival disappears) as $g_{1}/g_{2}$ approaches 1. The visibility vanishing at $g_{1}/g_{2}=0$ is trivial, since the unit cells are decoupled. However, the disappearance of the entanglement revival when $g_{1}/g_{2}\to 1$, which is the boundary of the topological phase, strongly suggests that the entanglement revival is closely connected with the edge mode.
Finally, while in this section we have exclusively discussed the logarithmic negativity, the universal revival can also be found in the mutual information (see A). The von Neumann and Rényi entropy densities exhibit a size-independent behavior as well (see B), suggesting that the topological properties of the system influence more generally the dynamical evolution and robustness of local correlations, both quantum and classical.
5 Conclusion
In this study of the fermionic and bosonic SSH models, characterized by the presence of a chiral symmetry, we explored the dynamics of entanglement of a single site with the rest of the chain, in the presence of decoherence which preserves the chiral symmetry. Our results reveal a remarkable entanglement revival which, for a fixed coupling ratio, acquires a universal profile when the system length exceeds the localization length of the edge states. Therefore, the effective length scale for the onset of universality can be identified with the localization length. These findings highlight a direct connection between the universal entanglement revival and the topological phase.
Acknowledgments
C.D. and M.-S.C. have been supported by the National Research Function (NRF) of Korea (Grant Nos. 2022M3H3A106307411 and 2023R1A2C1005588) and by the Ministry of Education through the BK21 Four program. S.C. acknowledges support from the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301602) and the National Science Association Funds (Grant No. U2230402).
Appendix A Mutual information
The mutual information is a measure of total correlations, both classical and quantum, between two subsystems. It provides a different perspective on the state of the system than entanglement measures, like the logarithmic negativity. Furthermore, the mutual information is usually easier to compute than most entanglement measures, making it a practical alternative for larger systems. Here, we show that the mutual information, like the logarithmic negativity, exhibits a universal entanglement revival, which indicates that universality is not limited to a specific entanglement measure.
The mutual information between subsystems $A$ and $B$ is defined as:
$\mathrm{I}_{\gamma}(A,B):=S_{\gamma}(\hat{\rho}_{A})+S_{\gamma}(\hat{\rho}_{B})-S_{\gamma}(\hat{\rho}),$ (22)
where $\hat{\rho}_{A/B}$ is the reduced density matrix of subsystem $A/B$ and $\hat{\rho}$ the full density matrix, while the Rényi entropy of order $\gamma$ is defined as:
In the limit $\gamma\rightarrow 1$, the Rényi entropy becomes the von Neumann entropy, $S_{\mathrm{vN}}(\hat{\rho})=-\mathrm{Tr}(\hat{\rho}\log_{2}\hat{\rho})$. However, computing the von Neumann entropy can still be computationally intensive for large systems. In the case with $\gamma=2$, the Rényi entropy simplifies to the second-order Rényi entropy, $S_{2}(\hat{\rho})=-\log_{2}\mathrm{Tr}(\hat{\rho}^{2})$, which is easier to compute and advantageous for large or complex systems.
Figure A1: Time dependence of the von Neumann mutual information in the fermionic model. $\mathrm{I}_{\mathrm{vN}}(\{2\},\{0,1,3,...\})$ indicates the von Neumann mutual information between the single-site subsystem $\{\hat{f}_{2}\}$ and the rest of the chain. In panels (a) and (b) we use $g_{1}/g_{2}=0.1$ and $0.5$, respectively. In both panels, $\Gamma/g_{2}=0.1$.
@@ Appendix A Mutual information
For fermionic Gaussian states, the von Neumann entropy can be calculated efficiently from the correlation function [57]. In Fig. A1, we can observe that the entanglement revival might disappear in certain limits, such as $g_{1}\ll g_{2}$, because the mutual information is not a proper measure of quantum entanglement. However, the universal behavior always persists, which is in agreement with the results for the logarithmic negativity presented in the main text. In addition, both measures suggest that, for a fixed coupling ratio $g_{1}/g_{2}$, the universal behavior can be observed when the size of the half-SSH model exceeds a certain crossover length. The consistency in behavior between the two measures underscores the robustness of the universal entanglement revival phenomenon, confirming once more that the universal behavior is determined by the inherent properties of the system.
@@
Figure A2: The revival behavior of von Neumann mutual information in the bosonic model for coupling ratios (a) $g_{1}/g_{2}=0.1$ and (b) $g_{1}/g_{2}=0.5$. Here, we set $\Gamma/g_{2}=0.1$.
@@
For the bosonic case, computing the von Neumann mutual information is more challenging than fermions, since the Hilbert space still grows exponentially with the chain length, increasing the difficulty in computing eigenvalues. Currently, we are able to calculate the von Neumann mutual information up to $L=13$. Figure A2 illustrates the entanglement dynamics for bosonic systems with different coupling ratios and chain lengths. As seen, despite the maximum length being much shorter than for fermions, the universal behavior is still present. Thus, the conclusions for bosons are very similar to those for fermions in Fig. A1, and to the behavior of the logarithmic negativity.
Due to the limitations in calculating the von Neumann mutual information for bosonic systems, we extended our analysis to the the Rényi mutual information, which can be obtained for larger system sizes (specifically, we consider $L=11,15,19$, and $23$). As shown in Fig. A3, also the Rényi mutual information displays a universal revival. The larger accessible system sizes allow us to obtain good convergence also in Fig. A3(b), where a larger coupling ratio $g_{1}/g_{2}=0.5$ is assumed. This extension further confirms that the universal behavior is not restricted to the von Neumann mutual information and logarithmic negativity, offering new evidence of its robustness in larger bosonic systems.
@@
Figure A3: Revival behavior of the Rényi mutual information in the bosonic model, for coupling ratios (a) $g_{1}/g_{2}=0.1$ and (b) $g_{1}/g_{2}=0.5$. We denote the Rényi mutual information between the single-site subsystem $\{\hat{f}_{2}\}$ and the rest of the chain by $\mathrm{I}_{2}(\{2\},\{0,1,3,\cdots\})$. The inset of panel (b) is a zoom-in of the main plot. We have used $\Gamma/g_{2}=0.1$.

### v2

@@ III.2 Entanglement measures
Quantifying quantum entanglement Plenio and Virmani (2007) is vital in quantum information theory, yet is highly non-trivial to calculate, especially, for mixed states. Common entanglement measures include entanglement entropy Bennett et al. (1996), concurrence Wootters (1998), and logarithmic negativity (or the closely related negativity) Plenio (2005); Eisert and Plenio (1999). However, entanglement entropy is applicable only to pure states, while concurrence is practical only for two-qubit systems, whether in pure or mixed states. For mixed states of more general systems, the logarithmic negativity (or negativity) has been widely used as one of very few practically computable entanglement measures. Nevertheless, as it requires finding the eigenvalues of the density matrix after partial transposition, its computational cost increases quickly with system size. To avoid complications arising from partial transposition, quantum mutual information is often used as well. Although mutual information is not an entanglement measure in the strict mathematical sense, it still captures many important aspects of quantum correlations as well as classical correlations. Below, we will investigate the entanglement content of the mixed state (19) using the logarithmic negativity. However, as we illustrate in Appendix A, our universal features are observed in the mutual information as well.
The logarithmic negativity between two subsystems $\mathbb{A}$ and $\mathbb{B}$ is defined by Plenio (2005)
$\mathcal{E}(\mathbb{A},\mathbb{B}):=\log_{2}\left\|\hat{\rho}^{T_{\mathbb{A}}}\right\|=\log_{2}\left\|\hat{\rho}^{T_{\mathbb{B}}}\right\|,$ (27)
where $\hat{\rho}^{T_{\mathbb{A}}}(\hat{\rho}^{T_{\mathbb{B}}})$ is the partial transpose of the density matrix $\hat{\rho}$ with respect to subsystem $\mathbb{A}$ ($\mathbb{B}$), and $\|\cdot\|$ denotes the trace norm. As mentioned above and seen in the definition (27), the main obstacle in computing the logarithmic negativity for large systems comes from the partial transposition: If not for the partial transposition, in many cases the trace norm may be calculated without much difficulty, by using the symmetries of the system to decompose the large density matrix into smaller blocks. More specifically, in our model the total number of particles is conserved, i.e., $\hat{n}_{\mathbb{\mathbb{A}}}+\hat{n}_{\mathbb{\mathbb{B}}},\hat{\rho}]=0.$ On the other hand, the partial transpose, $\hat{\rho}^{T_{\mathbb{A}}}$ or $\hat{\rho}^{T_{\mathbb{B}}}$, violates this symmetry, and one has to deal with the Hilbert space as a whole. Actually, in our model the difference $\hat{n}_{\mathbb{A}}-\hat{n}_{\mathbb{B}}$ in the number of particles is conserved Qiu et al. (2022), i.e, $\hat{n}_{\mathbb{A}}-\hat{n}_{\mathbb{B}},\hat{\rho}^{T_{\mathbb{A}}}]=0,$ but the invariant blocks under this symmetry are still exponentially large.
Recently, a new type of logarithmic negativity based on a partial time-reversal transformation has been suggested for fermionic systems Shapourian et al. (2017); Shapourian and Ryu (2019a, b). Compared to the conventional (or bosonic) logarithmic negativity, this fermionic logarithmic negativity was shown to better capture the contributions to entanglement from Majorana fermions. From our point of view, however, the key advantage lies in the fact that the partial time-reversal transformation of a fermionic Gaussian state remains Gaussian, allowing the fermionic logarithmic negativity to be calculated using a $2L\times 2L$ matrix of generalized Green’s functions Shapourian et al. (2017); that is, with a polynomial amount of resources and time. In our case, as the pairing correlations are zero, the fermionic logarithmic negativity is obtained directly from the $L\times L$ matrix of Green’s functions, see Eq. (26), and we could study half-chain SSH chains up to $L\sim 10^{3}$ without difficulties. For the bosonic case, on the other hand, the calculations are limited to relatively short chains.
III.3 Fermionic vs bosonic SSH models
Both bosonic and fermionic systems in the half-chain SSH model exhibit a chiral symmetry, characterized by the existence of a unitary operator that anti-commutes with the Hamiltonian. This symmetry leads to a degenerate zero-energy subspace and a symmetric single-particle energy spectrum around the zero energy. Additionally, the zero-energy edge mode reflects the non-trivial topological character of the band structure.
For large bosonic systems, the size of the Hilbert space poses significant computational challenges. Certain techniques, such as block diagonalization of $\hat{\rho}^{T_{\mathbb{A}}}$, are employed to alleviate the computational burden. In contrast, fermionic systems can benefit from the fact that the relevant state is Gaussian, since the partial time-reversal transformation of Eq. (22) remains Gaussian.
In the next Section, we will present our results for both bosonic and fermionic systems in parallel, to highlight the common universal behavior despite the differences summarized above. We attribute the common features of entanglement dynamics to the chiral symmetry and topological properties, especially the zero-energy edge modes.
IV Results
We now present our findings, which provide deeper insights into the relationship between dynamics and topological properties. Most of our calculations have been done using the symbolic quantum simulation framework Q3Choi (2020).
Figure 3: Universal entanglement revival, obtained by computing the logarithmic negativity $\mathcal{E}_{L}(\{2\},\{0,1,3,\dots\})$ between the single-site subsystem $\{\hat{f}_{2}\}$ and the rest of the chain. Panels (a) and (b) consider the fermionic model at $g_{1}/g_{2}=0.1$ and $0.5$, respectively. Panels (c) and (d) are for the bosonic model, also using $g_{1}/g_{2}=0.1$ and 0.5. Other parameters are $g_{2}=1$ (arbitrary units) and $\Gamma/g_{2}=0.1$.
IV.1 Universal entanglement revival
Figure 3 summarizes our main results, concerning the time evolution of entanglement between the second even site and the rest of the system (the qualitative features do not depend on a specific partitioning, see below). Figure 3 (a) and (b) refer to the fermionic half-chain SSH chain, where we computed the fermionic logarithmic negativity for two values of $g_{1}$ deep in the topological phase ($g_{1}\ll g_{2}$). On the other hand, Fig. 3 (c) and (d) show the conventional logarithmic negativity of the bosonic model in the same phase. In the fermionic case, we could calculate the fermionic logarithmic negativity for systems as large as $L\sim 10^{3}$; three representative sizes are shown in Fig. 3 (a) and (b). In the bosonic case, the calculations were limited to relatively small values of $L$, because of the computational difficulties discussed in Section III; the results for $L=3,5,\dots,11$ are shown in Fig. 3 (c) and (d). In both cases, two distinctive features are clearly seen: First, the entanglement revives after an initial decrease. Considering the very fragile nature of quantum entanglement in the presence of decoherence, an entanglement revival is unusual and has attracted much interest Chen et al. (2024); Życzkowski et al. (2001); Aolita et al. (2015); Tanaś and Ficek (2004); Benatti and Floreanini (2006); Cunha (2007); Abdel-Aty and Yu (2008); Ficek and Tanaś (2006, 2008); Paz and Roncaglia (2008); Drumond and Cunha (2009); Das and Agarwal (2009); Orszag and Hernandez (2010); Nunavath et al. (2023); Lakhfif et al. (2022); Amico et al. (2018). Second, the entire temporal profile of the entanglement has a universal dependence, i.e., it is almost unaffected by the system size. Although we only show the entanglement between site $\hat{f}_{2}$ and the remaining sites $\{\hat{f}_{0},\hat{f}_{1},\hat{f}_{3},\dots\hat{f}_{L}\}$ in Fig. 3, a universal revival behavior can be observed regardless of the specific choice of partitioning, as long as one subsystem is confined to the neighbor of the left edge. However, the visibility of the revival does depend on the partitioning.
As demonstrated in a recent work Chen et al. (2024), the entanglement revival itself is due to a self-purification process directly related to the chiral symmetry and the associated zero-energy mode (not necessarily localized to the edges), which behaves as a dark state over the system’s evolution. The system starts from a pure state with relatively large entanglement and, as it undergoes the decoherence process, all the eigenmodes $\hat{c}_{k\neq 0}$ of Eq. (11) (which have nonzero amplitudes on the odd sites $\hat{f}_{2i-1}$) are subject to decay and decoherence. Consequently, the system evolves into a mixture of various intermediate states, resulting in a reduced purity and the suppression of entanglement. After this intermediate stage, however, the system begins to recover its purity (hence, its entanglement), because all particles have the fate of decaying out, except for those occupying the zero-energy mode, which has nonzero amplitudes only on the even sites. Eventually, the system stabilizes to the final state, i.e., the zero-energy mode, which is immune to the decay process. The entanglement in the steady state is always determined by this zero-energy mode.
While the above arguments explain the presence of entanglement revival, its universal character is more intriguing, and is the main topic of the rest of the paper. In the following, we perform a detailed analysis of the size dependence, revealing that the universal behavior is attributed to the topological feature of the system, especially the zero-energy edge mode.
IV.2 Topological effects through the edge mode
In Fig. 3, we have seen that the entanglement revival deep in the topological phase ($g_{1}/g_{2}\ll 1$) is universal, having a dynamical profile independent of system size. On the other hand, as the boundary between the topological and trivial phases is approached ($g_{1}/g_{2}\to 1$), the evolution of entanglement shows significant variations especially for smaller systems. When $g_{1}/g_{2}=0.9$, for example, the dynamical profile for $L=3$ is distinct from the profiles of larger systems ($L=19,1001$), as shown in the inset of Fig. 4. We find that the dynamical profiles remain universal as long as the system size is sufficiently large compared with a certain length scale, directly related to the localization length $\xi$ of the zero-energy edge mode; see Eq. (1). In the following, we examine this crossover scale in detail.
We define two length scales concerning the entanglement revival dynamics: The crossover length $L_{\mathrm{cr}}$ is defined as the minimum system size which, at a given ratio $g_{1}/g_{2}$, achieves a universal time dependence. In other words, the dynamical profiles of systems larger than $L_{\mathrm{cr}}$ are all universal (within a certain numerical tolerance $\epsilon$). More specifically, let $\mathcal{E}_{L}(t)$ be the logarithmic negativity for a system of length $L$. Then, choosing a sufficiently large time $T$, at which the entanglement has reached its asymptotic value, we require that for all $L\geq L_{\mathrm{cr}}$
$\frac{1}{T}\int_{0}^{T}{dt}\,\left|\mathcal{E}_{L}(t)-\mathcal{E}_{\infty}(t)\right|\leq\epsilon\,\mathcal{E}_{\infty}(T).$ (28)
Numerically, we approximate the $L\to\infty$ limit $\mathcal{E}_{\infty}$ by taking $L=1001$. We also choose $T=20$ and $\epsilon=0.05$. Figure 4 plots the crossover length $L_{\mathrm{cr}}$ (blue empty circles) as a function of $g_{1}/g_{2}$. One can see that $L_{\mathrm{cr}}$ closely resembles the localization length $\xi$ of the edge mode (black solid line). For instance, when $g_{1}/g_{2}=0.9$, the crossover length matches the localization length $L_{\mathrm{cr}}\approx\xi\approx 19$. This value is highlighted by a horizontal dashed line in the main panel of Fig. 4. In the inset, the detailed entanglement profile with $L=19$ follows quite faithfully the universal dynamics at $L=1001$.
Figure 4: Comparison of the crossover length $L_{\mathrm{cr}}$ (blue empty circles) and the effective length $L_{\mathrm{eff}}$ (red filled circles) with the localization length $\xi$ of the edge mode (black solid line), as a function of the coupling ratio $g_{1}/g_{2}$. The red dashed line marks the values $L_{\mathrm{cr}}\approx 19$, computed at $g_{1}/g_{2}=0.9$. Inset: time dependence of the logarithmic negativity between the single-site subsystem $\{\hat{f}_{2}\}$ and the rest of the chain. Here we take $g_{1}/g_{2}=0.9$ and $L=3$, 1919, $1001$ (from top to bottom).
The second length scale we investigate is the effective length $L_{\mathrm{eff}}$ of the system. To define $L_{\mathrm{eff}}$, we consider the reduced density matrix $\hat{\rho}_{j}(t):=\operatorname{{Tr}}_{i=j+1}^{L}\hat{\rho}(t),$ obtained by tracing out sites $j+1,\cdots,L$. In general, the reduced density matrix $\hat{\rho}_{j}(t)$ represents a physically different state from the full density matrix $\hat{\rho}(t)$. However, in our case, it turns out that $\hat{\rho}_{j}(t)$ gives physically equivalent results to $\hat{\rho}(t)$, as long as $j$ is larger than a certain length scale, which we refer to as the effective length. That is, the effective length $L_{\mathrm{eff}}$ is the minimum length at which the reduced density matrix has the same properties of the full density matrix, $\hat{\rho}_{j}(t)\approx\hat{\rho}(t)$ for $j\geq L_{\mathrm{eff}}$. The comparison between $\hat{\rho}_{j}(t)$ and $\hat{\rho}(t)$ is done in terms of the logarithmic negativity, in a manner analogous to Eq. (28). More precisely, in Eq. (28) we substitute $\mathcal{E}_{L}(t)\to\mathcal{E}_{j}(t)$, where $\mathcal{E}_{j}(t)$ is the logarithmic negativity obtained from $\hat{\rho}_{j}(t)$. Ideally, we would compute $\hat{\rho}_{j}(t)$ from the full density matrix $\hat{\rho}(t)$ of the infinite half-chain SSH chain, i.e., in the limit $L\to\infty$. In practice, we can choose a sufficiently large value of $L$ and, as before, we have used $L=1001$ in our numerical calculations.
The explicit dependence of $L_{\mathrm{eff}}$ on $g_{1}/g_{2}$ is shown in Fig. 4 (red filled circles). It is seen that like the crossover length $L_{\mathrm{cr}}$, the effective length $L_{\mathrm{eff}}$ also agrees well with the localization length $\xi$ of the edge mode. In short, both $L_{\mathrm{cr}}$ and $L_{\mathrm{eff}}$ follow closely the localization length $\xi$, and this strongly suggest that the spatial range of the edge mode is a critical factor in determining the onset of the universal entanglement revival behavior.
Another feature that supports the topological origin of the entanglement revival is the visibility, defined as the difference between the asymptotic value ($t\to\infty$) and the minimum of the logarithmic negativity. The details of the visibility differ between the fermionic and bosonic cases, and are affected by system parameters as well. However, as illustrated in Fig. 5 for a long fermionic chain, we find the following generic qualitative features: The visibility is zero at $g_{1}/g_{2}=0$ and first increases with $g_{1}/g_{2}$, reaching a maximum around $g_{1}/g_{2}\approx 0.7$. Then, it sharply drops to zero (i.e., the entanglement revival disappears) as $g_{1}/g_{2}$ approaches 1. At $g_{1}/g_{2}=0$, the visibility vanishes for a trivial reason, i.e., because the unit cells are decoupled. However, the disappearance of the entanglement revival when $g_{1}/g_{2}\to 1$, which is the border between the topological and trivial phases, strongly suggests that the entanglement revival is closely connected with the edge mode.
Finally, while we have exclusively discussed the logarithmic negativity in this section, the universal revival can also be found in the mutual information (see Appendix A). The von Neumann and Rényi entropy densities exhibit a size-independent behavior as well (see Appnendix B), suggesting that the topological properties of the system affect the dynamical evolution and robustness of local correlations, both quantum and classical, in a universal way.
V Conclusion
In this study of the fermionic and bosonic SSH models, characterized by the presence of a chiral symmetry, we have explored the dynamics of entanglement of a single site with the rest of the chain, in the presence of decoherence which preserves the chiral symmetry. Our results reveal a remarkable entanglement revival which, for a fixed coupling ratio, acquires a universal profile when the system length exceeds the localization length of the zero-energy edge modes. The effective length scale for the onset of universality is identified with the localization length of the zero-energy edge modes. By revealing a direct connection between the universal entanglement revival and the topological phase, these findings enhance our comprehension of the impact of topological features on entanglement dynamics and suggest promising applications in adaptable quantum information processing systems, including photonic quantum computers and semiconductor quantum dot arrays. These results not only highlight the influence of topological features on entanglement dynamics but also motivate further investigation into various other decoherence scenarios. In particular, the effect of colored noise with chiral symmetry, which is relevant in experimental situations, is another interesting open question for future studies.
Acknowledgments
C.D. acknowledges the support from the STU Scientific Research Initiation Grant (Grant No. NTF25003T). C.D. and M.-S.C. have been supported by the National Research Function (NRF) of Korea (Grant Nos. 2022M3H3A106307411, 2023R1A2C1005588, RS-2024-00432563, RS-2024-00404854) and by the Ministry of Education through the BK21 Four program. S.C. acknowledges support from the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0301602) and the National Science Association Funds (Grant No. U2230402).
Appendix A Mutual information
Mutual information is a measure of total correlations, both classical and quantum, between two subsystems. It provides a different perspective on the state of the system than entanglement measures such as logarithmic negativity. Furthermore, mutual information is usually easier to compute than most entanglement measures, making it a practical alternative for larger systems. Here, we show that mutual information, like logarithmic negativity, exhibits a universal revival behavior, which indicates that universality is not limited to a specific entanglement measure.
The mutual information between subsystems $\mathbb{A}$ and $\mathbb{B}$ is defined by
@@ Appendix A Mutual information
where $\hat{\rho}_{\mathbb{A}/\mathbb{B}}$ is the reduced density matrix of subsystem $\mathbb{A}/\mathbb{B}$ and $\hat{\rho}$ the full density matrix, while the Rényi entropy of order $\gamma$ is defined as:
@@
For fermionic Gaussian states, the von Neumann mutual information can be calculated efficiently from the correlation function Casini and Huerta (2009). In Fig. A1, we observe that the revival behavior disappears in certain limits, such as $g_{1}\ll g_{2}$, because the mutual information is not a proper measure of quantum entanglement. However, the universal behavior always persists, which is in agreement with the results for the logarithmic negativity presented in the main text. In addition, both measures suggest that, for a fixed coupling ratio $g_{1}/g_{2}$, the universal behavior can be observed when the size of the half-chain SSH model exceeds a certain crossover length. The consistency in behavior between the two measures underscores the robustness of the universal entanglement revival phenomenon, confirming once more that the universal behavior is determined by the inherent properties of the system.
@@
Figure A2: The revival behavior of von Neumann mutual information in the bosonic model for coupling ratios for (a) $g_{1}/g_{2}=0.1$ and (b) $g_{1}/g_{2}=0.5$. In both panels, $\Gamma/g_{2}=0.1$.
@@
For the bosonic case, computing the von Neumann mutual information is challenging for large systems since the Hilbert space grows very fast with the system size and the relevant states are not Gaussian. Currently, we are able to calculate the von Neumann mutual information up to $L=13$. Figure A2 illustrates the entanglement dynamics for bosonic systems with different coupling ratios and chain lengths. As seen, despite the maximum length being much shorter than the fermion case, the universal behavior is still evident. Thus, the conclusions for bosons are very similar to those for fermions in Fig. A1, and to the behavior of the logarithmic negativity.
@@
Figure A3: Revival behavior of the Rényi mutual information of order 2 in the bosonic model for coupling ratios (a) $g_{1}/g_{2}=0.1$ and (b) $g_{1}/g_{2}=0.5$. Shown is the Rényi mutual information between site 2 and the rest of the chain. The inset of panel (b) is a zoom-in of the main plot. We have used $\Gamma/g_{2}=0.1$.
Due to the limitations in calculating the von Neumann mutual information for bosonic systems, we extended our analysis to the order-2 Rényi mutual information, which can be obtained for larger system sizes (specifically, we consider $L=11,15,19$, and $23$). As shown in Fig. A3, the Rényi mutual information also displays a universal revival. The larger accessible system sizes allow us to obtain good convergence also in Fig. A3(b), where a larger coupling ratio $g_{1}/g_{2}=0.5$ is assumed. This extension further confirms that the universal behavior is not restricted to the von Neumann mutual information and logarithmic negativity, offering new evidence of its robustness in larger bosonic systems.